# Adaptive Orchestration for Inference of Large Foundation Models at the Edge

### v1

@@ III-D Privacy and Security Considerations
IV Use Cases
IV-A Emergency Coordination in Smart Cities
@@
In a highly connected smart city environment, where autonomous AI agents are responsible for managing infrastructure, monitoring public safety, and responding to critical incidents, large-scale AI inference is crucial for maintaining operational efficiency. During emergency scenarios—such as large-scale blackouts, cyber-attacks on urban infrastructure, or natural disasters—the adaptive split inference system ensures real-time decision-making despite fluctuating resource availability.
@@
Consider a scenario where a massive earthquake disrupts transportation networks, damages critical infrastructure, and impairs traditional cloud connectivity. Smart city AI agents deployed across distributed MEC nodes play a pivotal role in orchestrating emergency response through AI Agents for Autonomous Coordination, specialized AI agents trained for disaster response, such as infrastructure monitoring bots, autonomous drones, and emergency service assistants, rely on continuous, high-throughput AI inference. For instance, these agents could process high-dimensional multi-modal data, including real-time video, LiDAR scans, and sensor data from IoT devices.
@@
In this environment, Adaptive Model Deployment are initially instantiated as Foundation Model partitions distributed across MEC nodes based on predefined computational capabilities and expected workloads. As infrastructure degradation leads to unstable connectivity and hardware failures, the system dynamically adjusts model partitions across available MEC nodes. For instance, if an AI agent controlling autonomous emergency drones detects a surge in demand for real-time object detection (e.g., identifying survivors in debris), the system triggers Model Re-Splitting (MR) and Split Revision (SR) to redistribute workloads efficiently.
When an MEC node reaches its computational threshold due to a high influx of emergency data streams, the Reconfiguration Broadcast (RB) mechanism ensures that AI agents can offload inference tasks to alternative nodes with idle capacity. The system dynamically revises the model split $S=(S1,S2,S3)$ into a more optimized configuration $S^{*}=(S^{*}_{1},S^{*}_{2},S^{*}_{3})$, continuously adapting to the dynamic conditions to maintain operational efficiency and robust performance.
IV-B Industry 4.0 Manufacturing Lines
@@
Modern manufacturing floors increasingly integrate edge AI for tasks like predictive maintenance, anomaly detection, and quality control, often under tight latency requirements. In such environments, multiple MEC nodes handle continuous data streams from high-speed sensors and robotic arms. When production ramps up unexpectedly, compute workloads spike and nodes near their capacity limits. The orchestrator responds by reassigning or even re-splitting the inference model across less-loaded nodes, preventing bottlenecks. Privacy-constrained segments, such as those inspecting proprietary designs, remain on trusted hardware, while more generic modules can be offloaded seamlessly to boost throughput.
IV-C Autonomous Vehicles and Intelligent Transport Systems
@@
Vehicle-to-infrastructure (V2I) services increasingly rely on advanced AI models for collision avoidance, route planning, and traffic flow optimization. Edge nodes at roadside units (RSUs) offer local compute to complement on-board vehicle processors, reducing latency while offloading computationally heavy layers. In busy urban corridors, traffic sensors and autonomous cars generate significant inference workloads. If congestion surges or a particular RSU becomes overloaded, the orchestrator redistributes model partitions among neighboring edge nodes, ensuring split inference scales effectively. Such real-time adaptivity allows vehicles to maintain continuous, low-latency awareness, meeting strict safety and efficiency standards critical in next-generation transport systems.
IV-D Expected Results
@@
TABLE I: Estimated values: Adaptive vs. Static Split Inference
Metric Static Adaptive
Latency 500-1000 ms 100-300 ms
Resource Utilization 50-60% 80-95%
Downtime incidents/hour) 5-10 0-2
QoS Compliance 60-70% 95-99%
@@
To quantitatively illustrate the operational improvements of the adaptive approach versus traditional static inference strategies, Table I provides comparative analysis based on previous results and estimates for a typical 5G-MEC environment [31, 32, 33, 34].
V Conclusions
This study has introduced an adaptive split inference orchestration framework designed to dynamically manage LFM partitions across heterogeneous edge nodes, addressing the inefficiencies inherent in static split inference methodologies. The framework establishes a foundation for real-time, QoS-aware, and privacy-preserving AI inference in edge computing environments, which is particularly crucial for latency-sensitive and resource-constrained applications. The proposed approach optimizes performance, enhances resource efficiency, and fortifies privacy preservation by leveraging real-time monitoring, workload redistribution, and dynamic reconfiguration.
The proposed orchestration model aligns with emerging objectives in the development of 6G networks by enabling intelligent, distributed AI processing at the edge. Its modular architecture facilitates seamless integration with existing edge orchestration systems while maintaining extensibility for future AI-driven optimizations. We argue that our proposed framework enhances system resilience and operational efficiency by dynamically adapting inference partitioning in response to fluctuations in network and compute conditions, both factors that will increase in complexity given future foundation model deployments in edge environments.
Proposed and necessary items for a research agenda on future work include:
@@ V Conclusions
Moreover, there are interesting investigation avenues around deployment strategies for mission-critical domains, such as autonomous systems, healthcare, and financial technology, to refine practical applicability. In addition, we should also pay attention to examining the broader impact of edge-driven AI inference on energy efficiency in smart city and industrial IoT applications to support sustainable AI deployment. Addressing these research challenges will be pivotal in advancing the state of distributed AI systems, particularly in the context of 6G-enabled networks and resource-constrained edge environments that drive AIaaS applications closer to the clients.
@@ Acknowledgment
This paper provides a scholarly description of the published Patent US20250071069A1 [35], along with references to published Patent US18/449811 [36], which have been conceived and authored while the authors were employed at International Business Machines, the current assignee of both patent applications.

### v2

@@ 3.4 Privacy and Security Considerations
4 Use Cases
4.1 Emergency Coordination in Smart Cities
@@
In a highly connected smart city environment, where autonomous AI agents are responsible for managing infrastructure, monitoring public safety, and responding to critical incidents [17], large-scale AI inference is crucial for maintaining operational efficiency. During emergency scenarios, such as regional blackouts, cyber-attacks on urban infrastructure, or natural disasters, adaptive split inference ensures real-time decision-making despite fluctuating resource availability.
@@
Consider a scenario where a massive earthquake disrupts transportation networks, damages critical infrastructure, and impairs traditional cloud connectivity. Smart city AI agents deployed across distributed MEC nodes can play a pivotal role in orchestrating emergency response through AI Agents for Autonomous Coordination, i.e. specialized AI agents trained for disaster response. This can be part of infrastructure monitoring bots, autonomous drones, and emergency service assistants, which rely on continuous, high-throughput AI inference. For instance, these agents need to process high-dimensional multi-modal data, including real-time video, LiDAR scans, and sensor data from IoT devices to provide continuous responses.
@@
In this environment, Adaptive Model Deployments are initially instantiated as foundation model partitions, which are distributed across MEC nodes based on predefined computational capabilities and expected workloads. As infrastructure degradation leads to unstable connectivity and hardware failures, the system dynamically adjusts model partitions across available MEC nodes. For instance, if an AI agent controlling autonomous emergency drones detects a surge in demand for real-time object detection (e.g., identifying survivors in debris), the system triggers split revision SR to redistribute workloads efficiently. Consequently, when an MEC node reaches its computational threshold due to a high influx of emergency data streams, the Reconfiguration Broadcast RB mechanism ensures that AI agents can offload inference tasks to alternative nodes with idle capacity. The system dynamically revises the model split $S=(S1,S2,S3)$ into a more optimized configuration $S^{*}=(S^{*}_{1},S^{*}_{2},S^{*}_{3})$, continuously adapting to the dynamic conditions to maintain operational efficiency and robust performance.
4.2 Industry 4.0 Manufacturing Lines
@@
Modern manufacturing floors increasingly integrate edge AI for tasks like predictive maintenance, anomaly detection, and quality control, often under tight latency requirements. In such environments, multiple MEC nodes (e.g., private 5G/6G-MEC base stations) may handle continuous data streams from high-speed sensors and robotic arms. When production ramps up unexpectedly, compute workloads spike and nodes near their capacity limits. The orchestrator then responds accordingly by reassigning or re-splitting the inference model across less-loaded nodes, preventing bottlenecks. Privacy-constrained segments, such as those inspecting proprietary designs, remain on trusted hardware, while more generic modules can be offloaded seamlessly to boost throughput.
4.3 Autonomous Vehicles and Intelligent Transport Systems
@@
Vehicle-to-infrastructure (V2I) services increasingly rely on advanced AI models for collision avoidance, route planning, and traffic flow optimization. Edge nodes at roadside units (RSUs) offer local compute to complement on-board vehicle processors, reducing latency while offloading computationally heavy layers. In busy urban corridors, traffic sensors and autonomous cars generate significant inference workloads. If congestion surges or a particular RSU becomes overloaded, the orchestrator redistributes model partitions among neighboring edge nodes, ensuring split inference scales effectively. Such real-time adaptivity allows vehicles to maintain continuous, low-latency awareness, meeting strict safety and efficiency standards that are critical in next-generation transport systems.
5 Expected Results and Quantitative Benefits
Table 4: Quantitative comparison of static and adaptive split inference across key performance dimensions in edge environments. Results are sourced from prior studies and testbed evaluations of LLM deployments over 5G-MEC infrastructures [43, 44, 45, 46, 47, 41, 28].
Performance Dimension Static Split Inference Adaptive Split Inference
Latency (End-to-End) High variance; $\sim$ 500-1000 ms typical in 5G-MEC scenarios. Cannot guarantee low latency under network fluctuation (may spike beyond 1 s). Low and stable; $\sim$ 100-300 ms under same conditions. Dynamically maintains latency below target (often 50%+ faster than static).
Throughput Limited by weakest link (device or network); lower overall. $<1\,\mathrm{req/s}$ on a baseline edge device alone. Higher via parallelism and load balancing; utilizes multiple nodes. Achieved $>5\,\mathrm{req/s}$ in adaptive multi-node setup ($\geq 5$× higher).
Resource Utilization $\sim$ 50-60% of available edge / cloud resources used. Static partition leads to idle resources on one side and potential overload on the other. $\sim$ 80-95% resource utilization. Orchestrator keeps both edge and cloud busy, scaling across nodes; more efficient use of CPU, GPU, and bandwidth.
QoS Compliance (SLA HitRate) Often poor under variability. QoS deadlines met in $\sim$ 60-70% of cases; frequent SLA violations when conditions deviate from design point. Near-guaranteed QoS even as conditions change. $\sim$ 95-99% of requests meet latency/SLA targets due to real-time adaptation (much fewer deadline misses).
Reliability & Failover Rigid deployment; single-point bottlenecks. $\sim$ 5-10 inference failures or timeouts per hour observed under load/network issues. Resilient, with dynamic re-routing and re-partitioning. Downtime incidents reduced to $\sim$ 0-2 per hour. Maintains service continuity by avoiding overloads.
Privacy Adherence Moderate. Fixed layer offload may send sensitive features off device. No ability to alter behavior for sensitive data—potential compliance issues. High. Can execute sensitive layers locally and limit data exposure. Adaptive policy balances performance with data confidentiality (privacy preserved without sacrificing QoS).
Table 4 presents a comparative analysis derived from prior studies, experimental testbeds, and empirical estimates for a representative 5G-MEC-enabled edge environment serving text-generation LLMs ranging from 7B to 13B parameters. For further technical depth and context, we refer to the comprehensive evaluations in [43, 44, 45, 46, 47, 41, 28].
In particular, it becomes evident that adaptive split inference offers significant improvements across multiple operational dimensions. First, it consistently achieves lower and more stable end-to-end latency, typically maintaining inference response times within 100–300ms, even under fluctuating 5G-MEC conditions, whereas static configurations frequently exceed 500ms with high variance. This translates to a latency reduction of over 50% in representative scenarios. Second, throughput is substantially improved: adaptive orchestration enables parallel execution across distributed nodes, reaching over 5 req/s compared to sub-1 req/s throughput observed in static, single-device deployments.
Resource utilization also benefits considerably. Static splits often lead to underutilized compute on one end (e.g., idle cloud GPUs) and overloading on the other, capping effective usage at 50–60%. In contrast, adaptive orchestration dynamically balances workloads, yielding sustained utilization rates of 80–95% across edge and cloud components. This directly contributes to higher QoS adherence: while static inference often fails to meet SLA constraints during network or load fluctuations, achieving target latency in only 60–70% of cases, adaptive inference maintains compliance for 95–99% of requests through real-time partition and placement adjustments.
Further, reliability under stress is markedly improved. Static deployments are prone to bottlenecks and failure cascades, with up to 10 inference errors or timeouts per hour observed under load. Adaptive strategies mitigate this via failover-aware routing and reconfiguration, reducing service interruptions to near-zero levels. In addition, privacy adherence is strengthened. While static splits may expose sensitive intermediate representations to untrusted infrastructure which can be targeted by malicious attackers, the adaptive approach can selectively retain privacy-critical layers on-device, enabling compliance with confidentiality requirements without degrading system performance.
Table 5 summarizes the discussed median figures reported across four public 5G–MEC studies [47, 41, 30, 48] and corresponding latency distributions are visualized in Figure 3 (static vs. adaptive CDF). The adaptive curve reaches the 95 % completion mark below $300$ ms, while the static curve stretches beyond $1$s, highlighting the practical QoS benefit of adaptive orchestration at runtime.
@@ 5 Expected Results and Quantitative Benefits
Table 5: Static vs. adaptive split inference for text-generation LLMs in typical 5G-MEC deployments (median values across the studies in [47, 41, 30, 48]).
Metric Static Split Adaptive Split
End-to-end latency 500–1000 ms 100–300 ms
Throughput (requests s-1) $\sim$1 $\sim$5
GPU/CPU utilisation 50–60 % 80–95 %
SLA hit-rate (400 ms budget) 60–70 % 95–99 %
Downtime incidents (per h) 5–10 0–2
@@
Across all evaluative dimensions (latency, throughput, utilisation, QoS robustness, and privacy), adaptive split inference consistently outperforms any static configuration. The small overhead of monitoring ($\leq$10 ms per cycle) and graph rewiring is amortised by hundreds of ms saved per request, yielding a net performance gain an order of magnitude larger than the orchestration cost. Consequently, adaptive split inference emerges as the decisive design choice for LLM deployment in heterogeneous, bandwidth-variable 5G/6G edge networks.
Figure 3: CDF of end-to-end inference latency for static (solid) vs. adaptive (dashed) split inference in a 5G-MEC scenario. 95 % of adaptive requests finish within 300 ms, while static requests may take up to 1s [47, 41, 30, 48].
6 Conclusions
This study has introduced an adaptive split inference orchestration framework designed to dynamically manage LFM partitions across heterogeneous edge nodes, addressing the inefficiencies inherent in static split inference methodologies. Our framework establishes a foundation for real-time, QoS-aware, and privacy-preserving AI inference in edge computing environments, which is particularly crucial for latency-sensitive and resource-constrained applications. The proposed approach optimizes performance, enhances resource efficiency, and fortifies privacy preservation by leveraging real-time monitoring, workload redistribution, and dynamic reconfiguration. In contrast to existing orchestration frameworks, we do not treat AI models as opaque containers, but instead support runtime model re-partitioning and privacy-aware layer placement, capabilities essential for deploying LFMs in dynamic edge environments. Our proposed orchestration model thus aligns with emerging objectives in the development of 6G networks by enabling intelligent, distributed AI processing at the edge. Its modular architecture facilitates integration with existing edge orchestrators while maintaining extensibility for future AI-driven optimizations.
To validate the practical relevance of our framework, we provided a comparative analysis grounded in prior studies, testbed results, and empirical evaluations. These results demonstrate consistent gains across latency, throughput, resource utilization, reliability, and privacy adherence, reinforcing the value of adaptive orchestration for real-world edge inference given the complexity of future foundation model deployments.
Nevertheless, several research directions emerge to further advance adaptive inference orchestration. Proposed and necessary items for a research agenda on future work include:
  1. 1.
@@ 6 Conclusions
Beyond technical extensions, domain-specific deployments in autonomous systems, healthcare, and industrial automation present high-impact opportunities for practical validation. Likewise, examining the role of edge-driven inference in reducing energy consumption and enabling sustainable AI operations across smart cities and IoT ecosystems opens new interdisciplinary research avenues. Addressing these challenges will be essential to realizing scalable, trustworthy, and privacy-preserving AIaaS architectures in the forthcoming 6G edge landscape.
@@ Acknowledgment
This paper provides a scholarly description of the published Patent US20250071069A1 [49], along with references to published Patent US18/449811 [50], which have been conceived and authored while the authors were employed at International Business Machines (IBM). IBM is the current assignee of both patent applications.